OpenStudy (anonymous):

Related Rates, Calculus: A ladder 10 feet long rests against a vertical wall. If the bottom of the ladder slides away from the wall at the rate of 1 feet/second, how fast is the top of the ladder sliding down the wall when the bottom of the ladder is 6 feet from the wall.

OpenStudy (zarkon):

use Pythagorean theorem

OpenStudy (anonymous):

it asks for a velocity, not a side of a triangle.

OpenStudy (anonymous):

it's asking for the rate of the one side of the triangle though

OpenStudy (anonymous):

set your first equation up like this x^2+y^2=10

OpenStudy (zarkon):

like I said...use \[x^2+y^2=z^2\]

OpenStudy (anonymous):

10^2 i mean

OpenStudy (anonymous):

no differentiate both sides in respects to t

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

then find out the height of y when x =6 and z = 10 to fnd the height and you'll have everything to put into the quation

OpenStudy (anonymous):

i got -.75 ft/sec but i haven't done this in like forever

OpenStudy (anonymous):